# Abdulrahman Alswaidan and Jeffrey D. Varner R.F. Smith School of Chemical and Biomolecular Engineering

### 06 Mar 2026 v1

@@ Acknowledgments
This work grew out of the CHEME-5820 Machine Learning and Artificial Intelligence Methods for Engineers course at Cornell University; the central idea arose while preparing a lecture on Boltzmann machines immediately after covering modern Hopfield networks. Financial market data used in Appendix G were obtained from Polygon.io.
@@ Appendix B Proof of Proposition 1
$\nabla^{2}E(\boldsymbol{\xi})=\mathbf{I}_{N}-\beta\,\mathbf{X}\,\mathbf{S}\bigl(\mathbf{p}(\boldsymbol{\xi})\bigr)\,\mathbf{X}^{\top}.$ (13)
@@
$\|\nabla^{2}E(\boldsymbol{\xi})\|_{\mathrm{op}}\;\leq\;\|\mathbf{I}_{N}\|_{\mathrm{op}}+\beta\,\|\mathbf{X}\,\mathbf{S}(\mathbf{p})\,\mathbf{X}^{\top}\|_{\mathrm{op}}\;\leq\;1+\beta\,\sigma_{\max}^{2}\,\|\mathbf{S}(\mathbf{p})\|_{\mathrm{op}},$ (14)
@@ Appendix C MALA Variant of Stochastic Attention
0: Memory matrix $\mathbf{X}\in\mathbb{R}^{N\times K}$, inverse temperature $\beta>0$, step size $\alpha\in(0,1)$, iterations $T$, initial state $\boldsymbol{\xi}_{0}\in\mathbb{R}^{N}$
@@
4:  $\boldsymbol{\epsilon}_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{N})$
@@
Compared with Algorithm 1, MALA replaces the single-line state update (line 4) with a propose-then-correct cycle (lines 2–13). The additional cost per step is one extra attention evaluation (line 6) and one energy evaluation (line 8), doubling the per-iteration $\mathcal{O}(NK)$ cost. In the limit $\alpha\to 0$ both algorithms coincide; at moderate $\alpha$ the accept/reject step removes $O(\alpha)$ discretization bias.
@@ Appendix D Load-Ratio Phase Diagram
Figure 3 showed the full phase diagram of attention concentration $C=1-H(\mathbf{a})/\log K$ over load ratio $K/d$ and inverse temperature $\beta$, as described in the main text.
@@ Appendix E Multi-Digit MNIST Generalization
To verify that the results reported for digit “3” in the main text are not specific to a single morphological class, we repeated the identical multi-chain protocol (30 chains, T=5, [^0] T{=}5{,}000, burn-in 2,0002{,}000, thin every 100 then sub-sample 5, $\beta{=}2000$, $\alpha{=}0.01$, K=100K{=}100 stored patterns per digit) on two additional MNIST digits chosen for their distinct morphological properties: digit “1” (simple stroke, near-degenerate orientation) and digit “8” (complex topology with two enclosed loops and high intra-class variance). All seven baselines (including the VAE) and the MALA variant were run with the identical protocol; results appear in Tables 2 and 3.
@@
Figure 4 and Figure 5 showed representative $4{\times}4$ sample grids for each method. The qualitative pattern was the same as for digit “3”: bootstrap outputs were exact stored copies, Gaussian perturbation was visually indistinguishable from bootstrap, random convex combinations produced uniform blurry averages, and both Langevin methods generated diverse, structured digits.
@@
The baseline comparison above fixes $\beta{=}2000$, which places the sampler deep in the retrieval regime: each chain stays near its seed pattern and the “generation” is local variation within a single energy basin. To show the full retrieval-generation trade-off on real images, Figure 6 displays $4{\times}4$ SA grids at four inverse temperatures alongside 16 stored patterns for reference. Table 4 reported the corresponding metrics.
Table 4: Temperature spectrum for digit “8” (K=100K{=}100, d=784d{=}784, $\alpha{=}0.01$, 30 chains). $\mathrm{SNR}=\sqrt{\alpha\beta/2d}$ is the per-step signal-to-noise ratio; the synthetic transition occurs near $\mathrm{SNR}\approx 0.02$$0.03$. As SNR falls through this band the sampler crosses from structured retrieval into diffuse generation. Values are mean $\pm$ SE across 30 chains.
@@ Appendix F Single-Chain Diversity Analysis
Table 5: Single-chain vs. multi-chain diversity on digit “3” (K=100K{=}100, d=784d{=}784, $\alpha{=}0.01$). $\mathrm{SNR}=\sqrt{\alpha\beta/2d}$. The multi-chain row reproduces the SA entry from Table 1. The diversity gap of $0.318$ at $\mathrm{SNR}{=}0.113$ is the initialization contribution; single-chain diversity exceeds the multi-chain value once SNR falls toward the transition band (0.020.020.030.03).
@@
Appendix G Financial Log-Return Generation
@@
G.1 Data and Memory Construction
@@
G.2 I.I.D. Pool: Marginal and Cross-Sectional Fidelity
@@
G.3 Sequential Generation and Temporal Properties
@@
Figure 11 (top row) showed the autocorrelation function of raw returns for five representative tickers. The generated ACF (blue) lay within the 99% confidence interval for white noise at all lags $\geq 1$, matching the historical pattern (red).
@@
Figure 11 (bottom row) showed the autocorrelation function of squared returns. The historical series (red) exhibits the characteristic slow-decaying positive autocorrelation of volatility clustering. The generated series (blue), however, showed ACF($g^{2}$) $\approx 0$ at all lags; the sampler did not reproduce this effect. We verified that at the main operating point ($\beta{=}25$, $T_{\text{inner}}{=}2{,}000$), ACF($g^{2}$) also collapses to ${\approx}0$ (mean over $424$ tickers: $0.0005$), confirming this is a property of equilibrium sampling and not an artifact of the choice of $\beta$.
@@
Table 6 made three theoretical predictions concrete. First, the novelty–fidelity trade-off is real and quantified: SA generates regime interpolations absent from the historical record ($\mathcal{N}=0.768\pm 0.001$) while bootstrap, which replays history verbatim, achieves $\mathcal{N}=0.000$; marginal fidelity runs in the opposite direction ($64.2\%$ vs. $99.3\%$ KS pass). Second, the Hopfield energy captures cross-sectional dependence independently of marginals: Frobenius error $26.3\%$ is unchanged by per-ticker affine correction, confirming that the energy geometry encodes correlation structure, not scale. Third, volatility clustering, a non-stationary phenomenon driven by regime shifts and leverage feedback [3], is exactly what a fixed-$\beta$ equilibrium sampler cannot reproduce; the absence is a theoretical prediction confirmed by experiment, not a tuning failure.
@@
Appendix H Simpsons Character Face Generation

### 10 Mar 2026 v2

@@ Acknowledgments
This work grew out of the CHEME-5820 Machine Learning and Artificial Intelligence Methods for Engineers course at Cornell University; the central idea arose while preparing a lecture on Boltzmann machines immediately after covering modern Hopfield networks. Financial market data used in Appendix H were obtained from Polygon.io.
@@ Appendix B Proof of Proposition 1
$\nabla^{2}E(\boldsymbol{\xi})=\mathbf{I}_{d}-\beta\,\mathbf{X}\,\mathbf{S}\bigl(\mathbf{p}(\boldsymbol{\xi})\bigr)\,\mathbf{X}^{\top}.$ (17)
@@
$\|\nabla^{2}E(\boldsymbol{\xi})\|_{\mathrm{op}}\;\leq\;\|\mathbf{I}_{d}\|_{\mathrm{op}}+\beta\,\|\mathbf{X}\,\mathbf{S}(\mathbf{p})\,\mathbf{X}^{\top}\|_{\mathrm{op}}\;\leq\;1+\beta\,\sigma_{\max}^{2}\,\|\mathbf{S}(\mathbf{p})\|_{\mathrm{op}},$ (18)
@@ Appendix C MALA Variant of Stochastic Attention
0: Memory matrix $\mathbf{X}\in\mathbb{R}^{d\times K}$, inverse temperature $\beta>0$, step size $\alpha\in(0,1)$, iterations $T$, initial state $\boldsymbol{\xi}_{0}\in\mathbb{R}^{d}$
@@
4:  $\boldsymbol{\epsilon}_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{d})$
@@
Compared with Algorithm 1, MALA replaces the single-line state update (line 4) with a propose-then-correct cycle (lines 2–13). The additional cost per step is one extra attention evaluation (line 6) and one energy evaluation (line 8), doubling the per-iteration $\mathcal{O}(dK)$ cost. In the limit $\alpha\to 0$ both algorithms coincide; at moderate $\alpha$ the accept/reject step removes $O(\alpha)$ discretization bias.
@@ Appendix D Load-Ratio Phase Diagram
Figure 3 shows the full phase diagram of attention concentration $C=1-H(\mathbf{a})/\log K$ over load ratio $K/d$ and inverse temperature $\beta$, as described in the main text.
@@ Appendix E Multi-Digit MNIST Generalization
To verify that the results reported for digit “3” in the main text are not specific to a single morphological class, we repeated the identical multi-chain protocol (30 chains, T=5, [^0] T{=}5{,}000, burn-in 2,0002{,}000, thin every 100 then sub-sample 5, $\beta{=}2000$, $\alpha{=}0.01$, K=100K{=}100 stored patterns per digit) on two additional MNIST digits chosen for their distinct morphological properties: digit “1” (simple stroke, near-degenerate orientation) and digit “8” (complex topology with two enclosed loops and high intra-class variance). All seven baselines (including the VAE) and the MALA variant were run with the identical protocol; results appear in Tables 3 and 4.
@@
Figures 4 and 5 show representative $4{\times}4$ sample grids for each method. The qualitative pattern was the same as for digit “3”: bootstrap outputs were exact stored copies, Gaussian perturbation was visually indistinguishable from bootstrap, random convex combinations produced uniform blurry averages, and both Langevin methods generated diverse, structured digits.
@@
The baseline comparison above fixes $\beta{=}2000$, which places the sampler deep in the retrieval regime: each chain stays near its seed pattern and the “generation” is local variation within a single energy basin. To show the full retrieval-generation trade-off on real images, Figure 6 displays $4{\times}4$ SA grids at four inverse temperatures alongside 16 stored patterns for reference. Table 5 reports the corresponding metrics.
Table 5: Temperature spectrum for digit “8” (K=100K{=}100, d=784d{=}784, $\alpha{=}0.01$, 30 chains). $\mathrm{SNR}=\sqrt{\alpha\beta/2d}$ is the per-step signal-to-noise ratio; the transition occurs near $\mathrm{SNR}^{*}=\sqrt{\alpha/(2\sqrt{d})}$ (Proposition 3); at d=784d{=}784 and $\alpha{=}0.01$ this gives $\mathrm{SNR}^{*}\approx 0.013$. As SNR falls through this region the sampler crosses from structured retrieval into diffuse generation. Values are mean $\pm$ SE across 30 chains.
@@ Appendix F Single-Chain Diversity Analysis
Table 6: Single-chain vs. multi-chain diversity on digit “3” (K=100K{=}100, d=784d{=}784, $\alpha{=}0.01$). $\mathrm{SNR}=\sqrt{\alpha\beta/2d}$. The multi-chain row reproduces the SA entry from Table 1. The diversity gap of $0.318$ at $\mathrm{SNR}{=}0.113$ is the initialization contribution; single-chain diversity exceeds the multi-chain value once SNR falls toward the entropy inflection point (Proposition 3).
@@
Mixing time.
At $\beta{=}2000$ the single chain remained in the basin of its seed pattern for the entire 50,00050{,}000-step run (${\approx}12$ s wall-clock on a laptop CPU): the max-cosine to the seed pattern stayed above $0.84$ at every thinned sample, and the chain never visited the basin of any other stored pattern. The inter-basin mixing time at this temperature therefore exceeds 50,00050{,}000 steps, consistent with the exponential-barrier scaling discussed in Section 4. Intra-basin mixing, by contrast, is fast: energy and max-cosine stabilize within the 2,0002{,}000-step burn-in, and consecutive thinned samples (every $100$ steps) have pairwise cosine distance $0.11\pm 0.01$, indicating decorrelation within $O(100)$ steps. At $\beta{=}200$ the picture changes: the chain crosses into a different basin within the first ${\approx}5{,}000$ steps and visits multiple basins over the 50,00050{,}000-step run, yielding the observed single-chain diversity of $0.796$.
Appendix G Protein Sequence Generation
To test stochastic attention on a non-image domain with discrete structure, we applied the full experimental pipeline to protein sequences from the Pfam RRM family (PF00076, RNA Recognition Motif) [24].
Data and encoding.
We downloaded the Pfam seed alignment (Stockholm format), removed columns with $50\%$ gaps and sequences with $30\%$ gaps, yielding K=68K{=}68 aligned sequences of length L=71L{=}71 positions over the 20 standard amino acids. Each sequence was one-hot encoded into $\mathbb{R}^{1420}$ ($20\times 71$), projected via PCA retaining 95% of variance ($1420\to 59$ dimensions), and $\ell_{2}$-normalized to form the memory matrix $\hat{\mathbf{X}}\in\mathbb{R}^{59\times 68}$.
Phase transition.
The entropy inflection analysis (Proposition 3) yielded $\beta^{*}\approx 3.85$ ($\mathrm{SNR}^{*}=0.018$), compared to the theoretical prediction $\beta^{*}=\sqrt{d}=7.68$ for random unit-norm patterns. The ratio of $0.50$ reflects the structured (non-random) geometry of the protein family: conserved positions reduce the effective variance of the similarities $e_{k}$, shifting the transition to lower $\beta$. We set $\beta_{\mathrm{ret}}=20\beta^{*}\approx 77$ ($\mathrm{SNR}=0.081$) for the structured-retrieval regime and $\beta_{\mathrm{gen}}=2\beta^{*}\approx 8$ ($\mathrm{SNR}=0.026$, just above the inflection) for the generation regime.
Protocol.
We ran the identical multi-chain protocol used for MNIST: 30 chains, T=5, [^0] T{=}5{,}000 iterations, burn-in 2,0002{,}000, thinning every 100 steps, 5 samples per chain, yielding S=150S{=}150 generated sequences. SA was run at both $\beta_{\mathrm{ret}}{=}77$ (retrieval) and $\beta_{\mathrm{gen}}{=}8$ (generation) to demonstrate the temperature knob on a non-image domain. All seven baselines (bootstrap, Gaussian perturbation, random convex combination, GMM-PCA, VAE with latent dim 8, MALA) were run with the same protocol. Generated PCA-space vectors were decoded back to amino acid sequences by inverse-PCA followed by per-position argmax over the 20 amino acid channels.
Metrics.
In addition to the standard metrics (novelty $\mathcal{N}$, diversity $\bar{\mathcal{D}}$, energy $\bar{E}$), we report two protein-specific measures: sequence identity to the nearest stored member (max over stored sequences of the fraction of non-gap positions with matching amino acids) and amino acid composition KL divergence ($D_{\mathrm{KL}}(p_{\mathrm{stored}}\|q_{\mathrm{gen}})$ over the 20 amino acid frequencies).
Table 7: Protein sequence generation on Pfam PF00076 (RRM family, K=68K{=}68, L=71L{=}71, d=59d{=}59 after PCA). Two SA rows demonstrate the temperature knob: $\beta{=}77$ (retrieval, $20\beta^{*}$) and $\beta{=}8$ (generation, $2\beta^{*}$). SeqID: maximum sequence identity to any stored sequence. KL: amino acid composition divergence from stored sequences (lower is better). Energy is positive at $\beta{=}8$ because samples explore off the attractor manifold. MALA acceptance rate: 99.8%.
Method $\mathcal{N}$ $\uparrow$ $\bar{\mathcal{D}}$ $\uparrow$ $\bar{E}$ $\downarrow$ SeqID KL $\downarrow$
Bootstrap (replay) $0.000\pm 0.000$ $1.000\pm 0.012$ $0.500\pm 0.000$ $0.645\pm 0.004$ 0.133
Gaussian perturbation $0.008\pm 0.000$ $1.002\pm 0.007$ $0.492\pm 0.000$ $0.633\pm 0.006$ 0.132
Random convex combination $0.510\pm 0.009$ $994\pm 0.009 0.0$$068\pm 0.001$ $0.562\pm 0.000$ 2.091
GMM-PCA (r=50r{=}50, C=10C{=}10) $0.606\pm 0.009$ $1.000\pm 0.007$ $0.077\pm 0.009$ $0.540\pm 0.003$ 0.731
VAE (latent=8{=}8) $0.621\pm 0.007$ $0.834\pm 0.022$ $0.118\pm 0.007$ $0.532\pm 0.003$ 0.416
MALA ($\beta{=}77$) $0.249\pm 0.004$ $0.992\pm 0.007$ $0.115\pm 0.005$ $0.609\pm 0.012$ 0.112
SA ($\beta{=}77$, retrieval) $0.243\pm 0.004$ $0.991\pm 0.007$ $0.120\pm 0.006$ $0.616\pm 0.012$ 0.107
SA ($\beta{=}8$, generation) $\mathbf{0.623\pm 0.006}$ $\mathbf{1.001\pm 0.006}$ $3.08\pm 0.06^{\dagger}$ $\mathbf{0.538\pm 0.006}$ 0.060
Figure 8: Protein sequence generation on PF00076. Left: Per-position amino acid frequencies for stored sequences (top) and SA-generated sequences (bottom). SA preserves the conserved positions of the RRM family while introducing variation at non-conserved sites. Right: Distribution of sequence identity to the nearest stored member for SA and GMM-PCA. SA generates sequences that are ${\approx}62\%$ identical to their nearest stored neighbor, while GMM-PCA produces more distant sequences (${\approx}54\%$) with much higher compositional divergence (KL=0.731=0.731 vs. $0.107$).
Table 7 reveals the same two-regime structure observed on MNIST. At $\beta{=}77$ (retrieval), SA generates sequences ${\approx}62\%$ identical to their nearest stored member with the lowest KL divergence of any method at this temperature ($0.107$). SA and MALA are again indistinguishable (MALA acceptance $99.8\%$), consistent with the MNIST results. Random convex combinations catastrophically distort the composition (KL=2.091=2.091) because averaging one-hot vectors in PCA space does not preserve per-position amino acid structure.
At $\beta{=}8$ (generation, $2\beta^{*}$), SA matches the VAE on novelty ($0.623$ vs. $0.621$) and exceeds it on diversity ($1.001$ vs. $0.834$), while achieving dramatically lower compositional divergence: KL=0.060=0.060 vs. $0.416$ for the VAE, a $6.9{\times}$ improvement. This means SA-generated sequences at generation temperature have the closest amino acid composition to the real RRM family of any method tested, including all baselines and SA at retrieval temperature. The temperature knob works as designed: lowering $\beta$ from retrieval to generation trades sequence identity ($0.616\to 0.538$) for novelty ($0.243\to 0.623$) while preserving family-level compositional fidelity. Energy rises above zero (samples leave the attractor manifold), but the Langevin gradient keeps the chain closer to the memory geometry than any learned baseline.
The entropy inflection analysis validates at this new dimensionality: the empirical $\beta^{*}{=}3.85$ is approximately half the random-pattern prediction $\sqrt{d}{=}7.68$ (Fig. 9). This discrepancy is expected: RRM sequences share conserved residues at many positions, reducing the effective variance of the query-key similarities and shifting the phase transition to lower $\beta$. The inflection criterion (Eq. 12) automatically accounts for this structure, providing the correct operating point without manual tuning.
Figure 9: Attention entropy $H(\beta)/\log K$ for the Pfam RRM memory (K=68K{=}68, d=59d{=}59). The empirical inflection point $\beta^{*}{=}3.8$ (dashed blue) occurs at roughly half the random-pattern prediction $\sqrt{d}{=}7.7$ (dotted green), reflecting the reduced effective similarity variance caused by conserved residues.
HMMER validation.
To provide a biologically meaningful validity metric beyond amino acid composition, we validated all generated sequences against the Pfam RRM profile HMM using HMMER 3.4 [25]. Gap characters were removed from each generated sequence before submission to hmmsearch. All $150$ generated sequences from every method pass at $E$-value <0.01<0.01, with SA at generation temperature achieving a median $E$-value of $8.1{\times}10^{-34}$ – far below the threshold and lower (better) than the stored sequences’ median of $2.5{\times}10^{-22}$. The generated sequences achieve better $E$-values because the generation process produces smoother, more canonical sequences that match the HMM consensus closely. The $100\%$ pass rate confirms that SA-generated sequences are biologically valid RRM family members, not arbitrary sequences that happen to share amino acid composition statistics.
Position-specific fidelity.
The global amino acid KL divergence aggregates frequencies across all positions. To verify that SA preserves position-specific conservation patterns, we computed the KL divergence independently at each of the L=71L{=}71 aligned positions and report the mean.
SA in the generation regime achieves a mean per-position KL of $2.92$, compared to $9.99$ for the VAE ($3.4{\times}$ lower), $7.52$ for bootstrap, and $9.09$ for GMM-PCA. This confirms that the advantage extends beyond global composition to position-specific conservation patterns: SA correctly identifies which positions are conserved and which are variable, preserving the position-specific amino acid distribution at each residue. MALA at generation temperature achieves a comparable $2.96$.
Pairwise coupling fidelity.
To address the concern that marginal frequency metrics ignore sequence-level dependencies, we computed the pairwise mutual information (MI) between residue positions. For a set of sequences, the MI between positions $i$ and $j$ is $\mathrm{MI}(i,j)=\sum_{a,b}p(a,b)\log\bigl[p(a,b)/(p(a)p(b))\bigr]$, where $p(a,b)$ is the joint frequency of amino acids $a$ and $b$ at positions $i$ and $j$. We computed MI matrices for each method’s generated sequences and measured the Pearson correlation with the stored family’s MI matrix.
SA at generation temperature achieves MI correlation r=0.871r{=}0.871, compared to r=0.525r{=}0.525 for the VAE – a $66\%$ improvement. Bootstrap replay achieves r=0.692r{=}0.692 and SA at retrieval temperature achieves r=0.733r{=}0.733. This demonstrates that SA-generated sequences reproduce the pairwise co-evolutionary coupling structure of the RRM family, capturing sequence-level dependencies that go beyond single-position statistics. The high MI correlation, combined with the low per-position KL and $100\%$ HMM pass rate, provides strong evidence that SA-generated sequences are structurally and functionally plausible RRM family members.
MALA at generation temperature.
SA and MALA produce near-identical results at the generation temperature ($\beta{=}8$): KL=0.060=0.060 vs. $0.058$, sequence identity $0.538$ vs. $0.537$, novelty $0.623$ vs. $0.616$, MI correlation $0.871$ vs. $0.887$ (MALA acceptance $99.8\%$). The ULA bias is negligible for protein sequences at both retrieval and generation temperatures.
Appendix H Financial Log-Return Generation
@@
H.1 Data and Memory Construction
@@
H.2 I.I.D. Pool: Marginal and Cross-Sectional Fidelity
@@
H.3 Sequential Generation and Temporal Properties
@@
Figure 13 (top row) shows the autocorrelation function of raw returns for five representative tickers. The generated ACF (blue) lay within the 99% confidence interval for white noise at all lags $\geq 1$, matching the historical pattern (red).
@@
Figure 13 (bottom row) shows the autocorrelation function of squared returns. The historical series (red) exhibits the characteristic slow-decaying positive autocorrelation of volatility clustering. The generated series (blue), however, shows ACF($g^{2}$) $\approx 0$ at all lags; the sampler did not reproduce this effect. We verified that at the main operating point ($\beta{=}25$, $T_{\text{inner}}{=}2{,}000$), ACF($g^{2}$) also collapses to ${\approx}0$ (mean over $424$ tickers: $0.0005$), confirming this is a property of equilibrium sampling and not an artifact of the choice of $\beta$.
@@
Table 8 makes three theoretical predictions concrete. First, the novelty–fidelity trade-off is real and quantified: SA generates regime interpolations absent from the historical record ($\mathcal{N}=0.768\pm 0.001$) while bootstrap, which replays history verbatim, achieves $\mathcal{N}=0.000$; marginal fidelity runs in the opposite direction ($64.2\%$ vs. $99.3\%$ KS pass). Second, the Hopfield energy captures cross-sectional dependence independently of marginals: Frobenius error $26.3\%$ is unchanged by per-ticker affine correction, confirming that the energy geometry encodes correlation structure, not scale. Third, volatility clustering, a non-stationary phenomenon driven by regime shifts and leverage feedback [4], is exactly what a fixed-$\beta$ equilibrium sampler cannot reproduce; the absence is a theoretical prediction confirmed by experiment, not a tuning failure.
@@
Appendix I Simpsons Character Face Generation